OpenStudy (anonymous):

A golf ball is hit into the air, and its height h in feet after t seconds is given by h(t)=-5t^2+130t. Determine the maximum height of the golf ball.

OpenStudy (anonymous):

find max height by finding axis of symmetry

OpenStudy (anonymous):

t = -b/2a b = 130 a = -5 t = -130/2(-5) = -130/-10 = 13 now solve for h(t) if t = 13